Prepare a Form 1040 and a Schedule A for Kofi

Kofi Allon, who is 32 years old and single, is employed as a technical consultant for a large electronics distributor. He lives at 678 Birch Street, LaMesa, CA 91941. Kofi's Social Security number is 412-34-5670 (date of birth 5/29/1987).


Kofi’s W-2 contained the following information:

 

Wages (box 1) =$ 75,350.62
Federal W/H (box 2) =$ 9,605.43
Social security wages (box 3) =$ 75,350.62
Social security W/H (box 4) =$ 4,671.74
Medicare wages (box 5) =$ 75,350.62
Medicare W/H (box 6) =$ 1,092.58

 

Kofi's other income includes interest on a money market account at the San Diego Credit Union of $275 and $265 in interest on a Certificate of Deposit at the Bank of California. He also made alimony payments of $650 per month to his ex-wife for ten months for a divorce agreement that was finalized on February 21, 2020. (Marta Tombe, Social Security number 412-34-5673).

 

In addition, he had the following other expenses for 2020:

 

Home mortgage interest$ 7,670
Auto loan interest1,325
Credit card interest730
Property taxes on personal residence2,342
State income taxes2,050
Unreimbursed hospital bills6,300
Doctor bills2,760
Other qualifying medical expenses725
Income tax preparation fee425
Job-hunting expenses925
Cash donations to his mosque (made ratably over the year and no single contribution was greater than $250)885

 

In may of 2020, Kofi’s house was broken into and his vintage 1961 Fender Stratocaster guitar was stolen. The guitar had a current fair market value of $15,750, and Kofi had paid $9,000 for the instrument in May of 2008. The guitar was uninsured.
Answer #1
In the books's of YY Ltd




Consolidated Statement of Comprehensive income as on 31st October, 2020



ParticularsYY LtdZZ LtdConsolidation

Sale£           150,000.00£ 84,000.00=150000+84000-3500-7000 =223500
LessCost of sale£              60,000.00£ 27,000.00=60000+27000-6000 = 81000

Gross Profit£              90,000.00£ 57,000.00£                                             142,500.00
LessOperating Exp.£              25,800.00£ 18,600.00=25800+18600 =44400

Profit befor Tax£              64,200.00£ 38,400.00£                                               98,100.00
LessTax£              12,000.00£   7,200.00=12000+7000= 19200

Profit for Tax£              52,200.00£ 31,200.00£                                               78,900.00






ParticularsAmt( In Pound)


Goods sold to ZZ Ltd£                7,000.00


Cost of Goods£                4,000.00


Profit£                3,000.00


Unsold Good's stock in ZZ Ltd£                7,000.00


Unrealised profit Eliminated=3000*85% = 2550


Unrealised profit Eliminated from NCI=3000*15% = 450


answered by: Mark